For the given sequence 1/n, 1/2n, 1/4n,…  Find the 16th term. Find sum of first sixteen terms where n is 156


1
Expert's answer
2021-07-18T17:12:39-0400

This is a geometric progression.

Let the 1st term be "a" and the common ratio "r".

Then "a=\frac{1}{n}" and "r=\frac{1}{2n} ÷ \frac{1}{n} = \frac{1}{2n} × \frac{n}{1} = \frac{1}{2}."

Let the nth term be "U_n." By the formula for the nth term of a geometric progression,

"U_n=ar^{n-1}"

So the 16th term of the geometric progression is

"U_{16}=ar^{15}=\frac{1}{n} × \left(\frac{1}{2}\right)^{15} \\=\frac{1}{n}×\frac{1}{32768} = \frac{1}{32768n}"


Let the sum of the first "n" terms of the sequence be "S_{16}." Then

"S_{16}=\dfrac{a\left(1-r^{n}\right)}{1-r}"

So the sum of the first 16 terms is

"S_{16}=\dfrac{a\left(1-r^{n}\right)}{1-r} \\
= \dfrac{1}{n}\dfrac{\left(1-\left(\dfrac{1}{2}\right)^{n}\right)}{1-\dfrac{1}{2}} \\"

"=\dfrac{1}{n}×2×\dfrac{65535}{65536} \\
=\dfrac{65535}{32768n}"

If "n = 156" then

"S_n=\dfrac{65535}{511808}"